OpenStudy (cold3693):

Find all the zeros of each function. f(x)=5x^3 + 6x^2 + x I know that you have to take out the x and that gives you 0. But what after?

OpenStudy (loser66):

\(5x^3+6x^2+x=x(5x^2+6x+1)\)

OpenStudy (loser66):

for some reason, It shows another number when I typed what I mean. :) the last one is ok. ( I can't delete the previous one)

OpenStudy (loser66):

now, by testing, you can see that if x = -1, then \(5x^2 +6x+1 =0\) hence, -1 is a root of the equation, make synthetic to get the leftover

OpenStudy (loser66):